OpenStudy (anonymous):

]How many liters of fluorine gas can react with 32.0 grams of sodium metal at standard temperature and pressure? Show all of the work used to find your answer. 2Na + F2 2NaF

sam (.sam.):

\[32~gNa \times \frac{1~molNa}{23~gNa} \times \frac{1~molF_2}{2~molNa} \times \frac{22.4~dm^3F_2}{molF_2}=?\]